An economy is forecasted to grow continuously so that the gross national product (GNP), measured in billions of dollars, after t years is given by:

GNP=60e0.03t

After how many years GNP is forecasted to reach RM88 billion?
1
Expert's answer
2020-12-17T04:43:51-0500

Solution to question150004


From the equation, "GNP=60e^{0.03t}"


Since we have the forecast value of GNP, we substitute it with RM88 billion as shown below


"RM88 billion=60e^{0.03t}"


"88=60e^{0.03t}, divide both sides by 60"


"\frac{88}{60}=e^{0.03t}"


"1.46667=e^{0.03t}"


"ln 1.46667=In (e^{0.03t}), introduce logarithmic"


"In \frac{1.46667}{0.03}= t"


"t=12.77 years"



GNP is forecast to reach RM 88 billion in 12.77 years